# When AI Evaluates Its Own Work: Validating Learner-Initiated, AI-Generated Physics Practice Problems

### v1

@@ III.3 Exit-survey feedback
The implemented system is unique in that it combines a traditional chatbot with an on-demand exercise generator, based on intent-detection (see Fig. 3). Across the comments, students agreed that the chatbot-problem-generation system offers a convenient single stop for quick topic overviews, formula lists, and immediately solvable practice tasks. Many praised the option to choose between conceptual multiple-choice and numerical calculation items, noting that the resulting questions often resembled past mini-exam formats and that the step-by-step solutions — when correct (llm-solution-is-correct) — served as clear worked examples (these statements correspond to includes-solution-strategy).
The quantitative ratings already hinted at a split verdict: roughly three-quarters of the 34 respondents judged the exercise generator “useful” or “rather useful,” yet only one in ten called it “very useful,” and 21 % found it “less useful.” When invited to justify these choices, students almost universally anchored their answers in two dimensions: pedagogical quality (difficulty, variety, accuracy) and workflow shortcomings (waiting time, presentation, language). Those who scored the system highly cited its ability to “produce a task on any topic I asked for” (corresponding to exercise-is-relevant-to-user-request), to supply the “most important formulas right away,” and to hand back step-by-step solutions that “look like mini-exam worked examples” (corresponding to llm-solution-is-correct and includes-solution-strategy). By contrast, lower ratings came from learners who felt the questions were “always the same type, just new numbers,” lacked diagrams or hints, or occasionally contained “wrong answer keys” (corresponding to llm-solution-is-correct), all of which undermined confidence and perceived value.
@@
The very same mechanics, however, generated the bulk of critical remarks. Roughly one-third of the cohort complained that tasks remained too basic or repetitive, providing practice in direct substitution rather than the “conceptual twists” typical of course assessments. Others met the inverse problem: single-step shortcuts or use of non-course notation (“another formula than the one we learned”) made some items feel deceptively hard or even wrong (corresponding to grounded-in-textbook). The absence of visuals (circuit diagrams, field sketches, etc.) was singled out in scores of comments; students trying to parse textual descriptions of spatial setups often had to “ask extra questions, losing minutes.” Errors in multiple-choice keys (corresponding to llm-solution-is-correct and single-correct-option) and “English instead of German even when I asked” (an issue that would fall under exercise-is-relevant-to-user-request, but which we failed to consider) further chipped away at trust. Finally, many disliked that the full solution appeared the moment they finished typing an answer: “I would rather reveal hints step-by-step or at least choose when to see the derivation; showing everything kills my own thinking process” (corresponding to solution-not-in-problem).
@@
In sum, the reasons students gave for their ratings converge on a clear message: the generator is already convenient and motivational, yet its educational impact remains capped by limited task diversity, occasional inaccuracies, missing visuals, and latency. Addressing these concrete deficiencies would include richer problem templates, stricter validation, optional hint-mode, embedded sketches, and faster response times (lower latency and higher throughput rate).
@@ IV Discussion
Putting it all together, we need to arrive at a pragmatic stack of metrics that is cost-efficient and latency-avoiding. Only three metrics emerge that can both be evaluated satisfactorily (Table 5) and strongly align with student preferences (Tables 68): includes-solution-strategy (an associated member of “Cluster A – Conceptual Depth” in Sect. II.4.2), task-is-specific-and-complete, and measurement-unit-is-clearly-stated (members of “Cluster B – Task Clarity & Feasibility”). Barely not making the somewhat arbitrary importance cut of $0.1$, but reliably measurable is llm-solution-is-correct (also in “Cluster B”).
Of high importance to learners are the two conceptual depth indicators difficulty (exercise-difficulty-measure) and Bloom level (bloom-level-of-exercise). Unfortunately, while establishing the ground truth, in a large number of cases, the expected-student-difficulty-level and expected-bloom-level could not be determined from the user prompt or chat context, and thus even if those two metrics could be judged reliably (which they cannot, see Table 5), frequently it would therefore still be impossible to judge if they match expectations. These measures, however, are closely correlated with includes-solution-strategy; the correlation is the result of the AI-based question generation, and we can only surmise how it came about. A possible explanation is that the model’s generation of a solution strategy is indicative of Chain-of-Thought reasoning [69], which in turn leads to greater “depth.”
@@
The metric grounded-in-textbook also had very few counter-examples; however, as the free–form feedback discussed in Sect. III.3 shows, if this occurs, it leads to confusion or irritation of the learners. The metric is a member of the isolated “Cluster C - Grounding & Relevance” in Sect. II.4.2. The strong correlation to exercise-is-relevant-to-user-request likely exists because if the question is not grounded in the textbook, it is likely way off from what the user asked for (such as the hallucinated questions about biology and the seasons).
@@ V Limitations
The evidence assembled here, while encouraging, must be interpreted with caution because several structural constraints narrow the study’s external validity. Our corpus of 543 generated tasks, though sufficient for a first pass at proxy discovery, remains too small and highly skewed to yield confirmatory statistics for roughly one third of the rubrics; a much larger, purposely balanced dataset will be required to stabilize minority classes and rule out ceiling artifacts. Equally important, every ground-truth label was provided by a single expert instructor. Without inter-rater checks, we cannot exclude systematic bias, particularly on inherently subjective constructs such as Bloom level or perceived difficulty, that may have propagated through both correlation graphs and random-forest models. We also worked with a convenience sample of just 34 volunteers drawn from one introductory physics cohort and tested them under time-limited laboratory conditions. Their click-through preferences may differ markedly from those of students in other institutions, disciplines, or self-paced home environments, limiting the generalizability of the choice-prediction layer.
A second cluster of limitations concerns instrumentation and causal inference. Our metric stack is text-only and therefore blind to diagrams or simulations that are central to many physics concepts; until multimodal scoring is integrated, overall task quality cannot be fully assessed. Moreover, the predictive models address which exercise a learner selects, not whether that choice enhances retention, transfer, or motivation over time; longitudinal outcome studies remain an open frontier. Because all feature-importance estimates are correlational, unobserved factors such as interface latency or learner fatigue could confound the apparent impact of surface cues like the presence of a solution strategy. Finally, the economic conclusions we draw — especially the cost-performance sweet spot of the o3-mini/low model — are contingent on a rapidly evolving marketplace; new reasoning engines or pricing schedules could upend the present optimization within months. Taken together, these constraints do not undermine the practical merits of our three-tier metric stack, but they do delineate a clear agenda for replication, multimodal expansion, and longitudinal validation.
@@ VI Outlook
The next step would be the implementation of a production version of the system in Fig. 3, widening and making more flexible the types of responses (beyond just numerical and 11-out-of-NN) and fully building the exercise validation step based on the insights gained.
Not yet under consideration is an instructional system that learns from the assessment outcomes and adapts continued instruction accordingly [45]. In other words, we have not endeavored to get to know the learner and, for example, ask appropriate follow-up questions. Instead, the assessment considered here was strictly on-demand, initiated by the learner; arguably, this is a serious shortcoming that would need to be addressed in future studies.
@@ VII Conclusions
Our results show that quality control for automatically generated physics problems can be achieved without an unwieldy battery of rubrics. A handful of indicators prove both measurable by affordable language-model calls and aligned with what students actually value. At the most basic level, every exercise must be internally consistent, unambiguous, and solvable with the information provided; tasks that fail any of these checks can be filtered out quickly and cheaply. Among the keepers, students gravitate toward problems that (i) offer a succinct roadmap or hint without revealing the answer, (ii) specify exactly how the response should be expressed (including the expected units) and (iii) signal an appropriate level of cognitive challenge. When those surface cues are present, solution correctness also rises, suggesting that clear structure and accurate content tend to co-occur.
Because many of the original rubrics overlap conceptually, a compact subset — roughly two checks for structural soundness, two for learner-visible appeal, and two lightweight keys for variety — delivers nearly the same predictive power as the full list while halving computational cost and latency. This streamlined stack therefore provides a pragmatic foundation for real-time formative assessment in physics: it is fast enough for conversational use, robust enough to keep low-quality items out of sight, and sensitive enough to present problems that students willingly engage with. Future work can build on this core, adding multimodal or adaptive layers, but the present study demonstrates that a small, well-chosen metric set is already sufficient to make on-demand AI-generated practice both trustworthy and useful.
@@ Acknowledgements.
We would like to thank the students who were part of our study, as well as Andreas Vaterlaus, who opened up his course for our study. We very much appreciated support by and discussions with Anna Kiepura, Jia Hong, Jessica Lam, and Richard Hahnloser. We would also like to thank Anna Kortemeyer for proofreading and critiquing our manuscript.

### v2

@@ III.3 Exit-survey feedback
The implemented system is unique in that it combines a traditional chatbot with an on-demand problem generator, based on intent-detection (see Fig. 4). Across the comments, students agreed that the chatbot-problem-generation system offers a convenient single stop for quick topic overviews, formula lists, and immediately solvable practice tasks. Many praised the option to choose between conceptual multiple-choice and numerical calculation items, noting that the resulting problems often resembled past mini-exam formats and that the step-by-step solutions — when correct (llm-solution-is-correct) — served as clear worked examples (these statements correspond to includes-solution-strategy).
The quantitative ratings already hinted at a split verdict: roughly three-quarters of the 34 respondents judged the problem generator “useful” or “rather useful,” yet only one in ten called it “very useful,” and one in five found it “less useful.” When invited to justify these choices, students almost universally anchored their answers in two dimensions: pedagogical quality (difficulty, variety, accuracy) and workflow shortcomings (waiting time, presentation, language). Those who scored the system highly cited its ability to “produce a task on any topic I asked for” (corresponding to exercise-is-relevant-to-user-request), to supply the “most important formulas right away,” and to hand back step-by-step solutions that “look like mini-exam worked examples” (corresponding to llm-solution-is-correct and includes-solution-strategy). By contrast, lower ratings came from learners who felt the problems were “always the same type, just new numbers,” lacked diagrams or hints, or occasionally contained “wrong answer keys” (corresponding to llm-solution-is-correct), all of which undermined confidence and perceived value.
@@
The very same mechanics, however, generated the bulk of critical remarks. Roughly one-third of the cohort complained that problem tasks remained too basic or repetitive, providing practice in direct substitution rather than the “conceptual twists” typical of course assessments. Figure 7 bears this out: in the majority of situations where the human expert was able to discern the expected difficulty, it was “intermediate,” while the majority of generated questions was only “foundational.” Others met the inverse problem: single-step shortcuts or use of non-course notation (“another formula than the one we learned”) made some items feel deceptively hard or even wrong (corresponding to grounded-in-textbook). The absence of visuals (circuit diagrams, field sketches, etc.) was singled out in scores of comments; students trying to parse textual descriptions of spatial setups often had to “ask extra questions, losing minutes.” Errors in multiple-choice keys (corresponding to llm-solution-is-correct and single-correct-option) and “English instead of German even when I asked” (an issue that would fall under exercise-is-relevant-to-user-request, but which we failed to consider) further chipped away at trust. Finally, many disliked that the full solution appeared the moment they finished typing an answer: “I would rather reveal hints step-by-step or at least choose when to see the derivation; showing everything kills my own thinking process” (corresponding to solution-not-in-problem).
@@
In sum, the reasons students gave for their ratings converge on a clear message: according to their feedback, the generator is already convenient and motivational, yet its educational impact remains capped by limited task diversity, occasional inaccuracies, missing visuals, and latency. Students generally wanted harder problems; they also wanted hints in case they get stuck, but expected them to only show up on-demand. Addressing these concrete deficiencies would include richer problem templates, stricter validation, optional hint-mode, embedded sketches, and faster response times (lower latency and higher throughput rate).
@@ IV Discussion
RQ1 is addressed by the model–ground-truth benchmarking in Table 5. RQ2 is addressed by the correlation analysis (Fig. 6) and random-forest importance rankings (Tables 68), together with Sect. III.3. RQ3 is answered in this section by proposing a compact metric stack based on the overlap of reliability, relevance, and automatic assessability.
Putting it all together, we need to arrive at a pragmatic stack of metrics that is cost-efficient and latency-avoiding. Only three metrics emerge that can both be evaluated satisfactorily (Table 5) and strongly align with student preferences (Tables 68): includes-solution-strategy (an associated member of “Cluster A – Conceptual Depth” in Sect. II.5.2), task-is-specific-and-complete, and measurement-unit-is-clearly-stated (members of “Cluster B – Task Clarity & Feasibility”). The latter two metrics combined are essentially, “do I know exactly what is asked for, and do I have everything needed to answer?”
Barely not making the somewhat arbitrary importance cut of $0.1$, but reliably measurable is llm-solution-is-correct (also in “Cluster B”). This metric generally cannot be the reason for students picking a problem, since they would not know if the solution is correct or not before attempting it, but it correlates with their choice. Anecdotal evidence shows that more modern LLMs than the ones tested here, namely models such as GPT-5 Thinking [78], solve about 90% of typical physics problems at the introductory level correctly [79], so today a simple algorithm of the LLM calculating the answer to a generated question and then comparing that to the generated solution would likely have much higher reliability than the one found in Table 5. In any case, in a practice problem, students would get very confused if the correct answer were not recognized as such.
Of high importance to learners when selecting a problem to work on are the two conceptual depth indicators difficulty (exercise-difficulty-level) and Bloom level (bloom-level-of-exercise). Unfortunately, while establishing the ground truth, in a large number of cases, the expected-student-difficulty-level and expected-bloom-level could not be determined from the user prompt or chat context, and thus even if those two metrics could be judged reliably (which they cannot, see Table 5), frequently it would therefore still be impossible to judge if they match expectations (this is distressing, since earlier studies of the quality of student-authored problems used Bloom as one of the main metrics [15]). These measures, however, are closely correlated with includes-solution-strategy; the correlation is the result of the AI-based problem generation, and we can only surmise how it came about. A possible explanation is that the model’s generation of a solution strategy is indicative of Chain-of-Thought reasoning [80], which in turn leads to greater “depth.”
@@
The metric grounded-in-textbook also had very few counter-examples; however, as the free–form feedback discussed in Sect. III.3 shows, if this occurs, it leads to confusion or irritation of the learners. The metric is a member of the isolated “Cluster C - Grounding & Relevance” in Sect. II.5.2. The strong correlation to exercise-is-relevant-to-user-request likely exists because if the problem is not grounded in the textbook, it is likely way off from what the user asked for (such as the hallucinated problems about biology and the seasons).
@@ V Limitations
The evidence assembled here, while encouraging, must be interpreted with caution because several structural constraints narrow the study’s external validity — it is thus exploratory in nature. Our corpus of 543 generated problems, though sufficient for a first pass at proxy discovery, remains too small and highly skewed to yield confirmatory statistics for roughly one third of the metrics; a much larger, purposely balanced dataset will be required to stabilize minority classes and rule out ceiling artifacts.
Equally important, every ground-truth label was provided by a single expert instructor, due to the high workload involved. Without inter-rater checks, we cannot exclude systematic bias, particularly on inherently subjective constructs such as Bloom level or perceived difficulty, that may have propagated through both correlation graphs and random-forest models.
We also worked with a convenience sample of just 34 volunteers drawn from one introductory physics cohort and tested them under time-limited laboratory conditions. Their click-through preferences between two candidate problems (captured by the chosen variable) may differ markedly from those of students in other institutions, disciplines, or self-paced home environments, limiting the generalizability with respect to student preferences. Moreover, we did not collect the corresponding instructor choices; consequently, the choice-prediction layer reflects learner-centered rather than instructor-centered notions of preference and relevance.
A second cluster of limitations concerns instrumentation and causal inference. Our metric stack is text-only and therefore blind to diagrams or simulations that are central to many physics concepts; until multimodal scoring is integrated, overall task quality cannot be fully assessed. Moreover, the predictive models address which problem a learner selects, not whether that choice is a true proxy for quality or enhances retention, transfer, or motivation over time; longitudinal outcome studies remain an open frontier. Because all feature-importance estimates are correlational, unobserved factors such as interface latency or learner fatigue could confound the apparent impact of surface cues like the presence of a solution strategy. Finally, the economic conclusions we draw — especially the cost-performance sweet spot of the o3-mini/low model — are contingent on a rapidly evolving marketplace; new reasoning engines or pricing schedules could upend the present optimization within months. Taken together, these constraints do not undermine the practical merits of our three-tier metric stack, but they do delineate a clear agenda for replication, multimodal expansion, and longitudinal validation.
Finally, because relevance is operationalized via student choice in a constrained exam-preparation workflow, the resulting importance rankings may overweight immediately approachable features (e.g., visible solution steps) and may not align with instructor goals such as coverage or desirable difficulty; ideally, future work should replicate the preference study with an instructor cohort and, ideally, relate both sets of preferences to learning outcomes.
@@ VI Outlook
The next step would be the implementation of a production version of the system in Fig. 4, widening and making more flexible the types of responses (beyond just numerical and 11-out-of-NN) and fully building the problem-validation step based on the insights gained. In production usage, a more diverse and balanced set of AI-generated problems could be obtained, and with more course instructors involved, at least for a subset of those problems, more than one expert could contribute ground truth. The system could thus be iteratively improved over several semesters based on a broader dataset and more broadly based ground truth.
Likely based on the training corpus of LLMs, the set of problems generated in this study was similar to standard end-of-the-chapter problems in older, drill-focussed textbooks [81]; it lacked variety, interest, and depth (as also reflected by the student survey). A production system may need to implement a flow of several LLM-calls that better simulate creativity. Not yet under consideration was an instructional system that learns from the assessment outcomes and adapts continued instruction accordingly [52]. In other words, we have not endeavored to get to know the learner and, for example, ask appropriate follow-up questions. Instead, the assessment considered here was strictly on-demand, initiated by the learner; arguably, this is a serious shortcoming that would need to be addressed in future studies.
@@ VII Conclusions
Our results show that the validation of automatically generated physics problems can be achieved without an unwieldy battery of metrics. A handful of indicators prove measurable and aligned with expert evaluation by affordable language-model calls and at the same time aligned with what students actually value — this is not necessarily what is sufficient, but what is feasible at the moment. The stack of metrics evaluates correctness, solvability, and alignment with features desired by students, not pedagogical quality (in fact, the metrics most closely associated with pedagogical quality, Bloom-level and difficulty, proved to be hard to assess by both human and AI). At the most basic level, every problem must be internally consistent, unambiguous, and solvable with the information provided; tasks that fail any of these checks can be filtered out quickly and cheaply. Among the keepers, students gravitate toward problems that (i) offer a succinct, on-demand roadmap or hint without revealing the answer, (ii) specify exactly how the response should be expressed (including the expected units) and (iii) signal an appropriate level of cognitive challenge. When those surface cues are present, solution correctness also rises, suggesting that clear structure and accurate content tend to co-occur.
Because many of the original metric items overlap conceptually, a compact subset — roughly two checks for structural soundness, two for learner-visible appeal, and two lightweight keys for variety — delivers nearly the same predictive power as the full list while halving computational cost and latency. This streamlined stack therefore provides a pragmatic foundation for real-time formative assessment in physics: it is fast enough for conversational use, robust enough to keep low-quality items out of sight, and sensitive enough to present problems that students willingly engage with.
@@ Acknowledgements.
We would like to thank the students who were part of our study, as well as Andreas Vaterlaus, who opened up his course for our study. We very much appreciate the support from and discussions with Jessica Lam, Anna Kiepura, and Richard Hahnloser. We would also like to thank Anna Kortemeyer for proofreading and critiquing our manuscript.